OpenStudy (anonymous):

Write the equation of the line in slope-intercept form that passes through the points (-2, 6) and (-1, 4).

OpenStudy (anonymous):

y = mx + b Find the slope first. Do you know how to do that?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

(y-y1)/(y2-y1)=(x-x1)/(x2-x1)

OpenStudy (anonymous):

m=y2-y1/x2-x1

OpenStudy (anonymous):

equ y-y1=m(x-x1)

OpenStudy (anonymous):

got it